OpenStudy (anonymous):

Ok.....here is another..arman, where ya at? lol :) An investment of 40,000 was made by a business club. The investment was split into three parts and lasted for one year. The first part of the investment earned 8% interest, the second 6%, and the third 9%. Total interest from the investments was 3060. The interest from the first investment was 2 times the interest from the second. Find the amounts of the three parts of the investment.

OpenStudy (anonymous):

i = prt

OpenStudy (anonymous):

Arman...was this one too much for you? Ha! :)

OpenStudy (anonymous):

Omg... give me a sec

OpenStudy (anonymous):

LOL!

OpenStudy (anonymous):

R U sure the next to last sentence is accurate?

OpenStudy (anonymous):

hmmm lemme c

OpenStudy (anonymous):

The interstate formula doesn't do anything, the interstate values are totals

OpenStudy (anonymous):

Yes it's right. @ Arman....what?!

OpenStudy (anonymous):

This problem is a owletgit lolol

OpenStudy (anonymous):

Yeah, I know it's tough~

OpenStudy (anonymous):

I think it is stated incorrectly. The next to last statement should be about the amounts. 3060 = .08(2x) + .06(x) + .09(40000 - 3x)

OpenStudy (anonymous):

lemme c

OpenStudy (anonymous):

306000 = 8(2x) + 6(x) + 9(40000 - 3x) 306000 = 16x + 6x + 360000 - 27x 306000 = -5x + 360000 5x = 54000 x = 10500

OpenStudy (anonymous):

8% = $21000 6% = $10500 9% = $8500

OpenStudy (anonymous):

nope. that's what it says.....

OpenStudy (anonymous):

do the first part is 21000?

OpenStudy (anonymous):

*so

OpenStudy (anonymous):

A is 3889.83, b is 2593.22 and c is 33516.95... the values are not 100% precise because a endless rationale was involved

OpenStudy (anonymous):

but the total was only like 3060.

OpenStudy (anonymous):

Total is 40k

OpenStudy (anonymous):

interest i mean

OpenStudy (anonymous):

I see what I did wrong. Tommy another sec

OpenStudy (anonymous):

let x, y and z be the 1st, 2nd and the 3rd investments respectively. according to the given informations, we have the following three equations; \[x+y+z=40,000 \rightarrow(1)\] \[0.08x+0.06y+0.09z=3060 \rightarrow(2)\] \[0.08x=2(0.06y) \rightarrow(3)\]

OpenStudy (anonymous):

.......

OpenStudy (anonymous):

solving three equations, we get that: first investment x=18,000 second investment y=12,000 third investment z=10,000

OpenStudy (anonymous):

the three*

OpenStudy (anonymous):

is something wrong?

OpenStudy (anonymous):

I don't know if that makes any sense to you. do you have the final answer anyway?

OpenStudy (anonymous):

Lol, ok. lemme c.

OpenStudy (anonymous):

Oh my gosh.........it's right!

OpenStudy (anonymous):

A is 20193.66 b is 6731.28 c is 13025.06

OpenStudy (anonymous):

Actually the 18,12, and 10 was right!

OpenStudy (anonymous):

^^ I knew it was right :P .. I was just asking if it makes sense to you? lol

OpenStudy (anonymous):

made*

OpenStudy (anonymous):

I got a problem for kelly: there is a integer, x, that is 10 digits long that if entered telephonically goes to you. What is x?

OpenStudy (anonymous):

lol.. this is a hard one for her :P

OpenStudy (anonymous):

arman, if you told me earlie, I would have sent you the answer and let you post it.. ;)

OpenStudy (anonymous):

anyway guys, it's my bed time already.. don't forget to click "become a fan", next to my name.. I hope you get the right answer arman!

OpenStudy (anonymous):

come on!

OpenStudy (anonymous):

I don't do "problems"!

OpenStudy (anonymous):

...........obviously.......lol. sigh.

OpenStudy (anonymous):

A telephone!

OpenStudy (anonymous):

A telephone... that's no number! And think of it this way, you can ask me all the math problems you want ;)

OpenStudy (anonymous):

lol...silly,

OpenStudy (anonymous):

you know you blew that last one! ;)

OpenStudy (anonymous):

Haha i think I'm a bit young anyways :P

OpenStudy (anonymous):

That was a bs question though -.-

OpenStudy (anonymous):

How is that?

OpenStudy (anonymous):

Cause nobody would several need to do a problem like that -.- and you knew the answer

OpenStudy (anonymous):

?

OpenStudy (anonymous):

"?"?

OpenStudy (anonymous):

Cause nobody would several need to do a problem like that -.- and you knew the answer

OpenStudy (anonymous):

whats that mean?

OpenStudy (anonymous):

Well you knew the answers lol and there is no real life problem where you know a total investment and interstate without knowing the values but you somehow know that one interstate is double the other? It's a useless trivia question... no offense :)

OpenStudy (anonymous):

I didn't know the answer. Trust me! That other guy knew it.

OpenStudy (anonymous):

Subjugation whateverrrr its still trivially useless :P

OpenStudy (anonymous):

Ahhhhhh* not Subjugation... stupid autospellcheck

OpenStudy (anonymous):

No I have to go grab a dictionary!

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

to make submissive or subservient; enslave.....what the hell?

OpenStudy (anonymous):

eh,

OpenStudy (anonymous):

What I don't get is why do you have all these words problems?

OpenStudy (anonymous):

math homework, duh!

OpenStudy (anonymous):

Would you prefer I give you regular problems?

OpenStudy (anonymous):

guess not...

OpenStudy (anonymous):

Lol whatsup grade are you in? And id love to help but I think the others got you covered-i have to put my brain to my work, I can chat though

OpenStudy (anonymous):

What... Damn spellcheck

OpenStudy (anonymous):

Or not...